OpenStudy (anonymous):

y varies directly as the square of x. When x = 5, y = 100. Find y when x = 6. y = 105 y = 120 y = 144 y = 150

Parth (parthkohli):

What's the square of \(x\) when \(x = 5\)?

OpenStudy (anonymous):

um...........idk

Parth (parthkohli):

What's the square of 5?

OpenStudy (anonymous):

25?

Parth (parthkohli):

Yeah, but let's leave that for a minute.\[y \text{ varies directly with }x^2 \implies y = kx^2\]

Parth (parthkohli):

We have to find \(k\), and we can do that by substituting \(x = 5\) and \(y = 100\).\[100 = k(5^2) \]\[100=25k\]\[ k = 4\]Does that make sense?

OpenStudy (anonymous):

yes

Parth (parthkohli):

Now you have to find \(y \) when \(x = 6\). Recall our equation\[ y = kx^2\]We already know that k = 4 and x = 6. Substituting,\[y = 4(6^2) \]\[ \implies 4(36)\]\[\implies\cdots \]

Parth (parthkohli):

What is 4(36)?

OpenStudy (anonymous):

144

Parth (parthkohli):

Right, and that is it.

OpenStudy (anonymous):

thanks!!